# Leveraging data-driven weather models for improving numerical weather prediction skill through large-scale spectral nudging

### v1

@@ 6.6 Computational cost of spectral nudging
The current implementation of spectral nudging in GEM with the presented optimal configuration leads to an increase of computational cost by approximately 25%. Due to the proof-of-concept nature of this study, no optimization of the informatics code has been performed. Therefore, the cost could be considerably reduced through future code optimization, especially for the spectral filter. Further improvements could be attained by considering time-varying nudging increments, where nudging would only be applied at time-steps close to the inference times of GC. These considerations however lie beyond the scope of this study.
@@ 7 Summary and Future Work
The recent emergence of AI models has begun to disrupt the operational weather forecasting paradigm which has long been dominated by the physics-based NWP models. Naturally, serious questions about the future role of AI and NWP models in operational forecasting are arising. For instance, are AI models capable of adding any reliable and meaningful value for weather services, or will NWP models eventually disappear from the forecasting landscape and be replaced by AI models? This study attempts to address these pressing questions by choosing the GEM and GC models to represent the two forecasting approaches.
@@
First, by comparing and contrasting the GEM and GC models, their relative strengths and weaknesses are identified. Skill verification against observations and GDPS analyses reveals that GC suffers from excessive fine-scale smoothing, affecting scales as large as 2500 km beyond day 3 of a 10-day forecast. However, despite smoothing, the results demonstrate that large-scale features (wavelength$>$2500 km) in GC exhibit better skill compared to GEM, which is evidenced by increased spectral coherence relative to the GDPS analyses and reduced STDE against radiosonde observations.
@@
Inspired by the improved large-scale skill of GC, a hybrid NWP-AI system – namely, GDPS-SN – has been developed to produce global forecasts where GEM-predicted large scales in horizontal wind and temperature are spectrally nudged toward inferences from GC. This hybrid system is capable of generating real-time forecasts with accuracy that surpasses ECCC’s operational GPDS. For example, RMSE of the 500-hPa geopotential height is reduced by 5-10%, with an overall predictability improvement of 12-18 hours, peaking at day 7 of a 10-day forecast. Notably, trajectories of tropical cyclones are predicted with enhanced accuracy without significant changes in intensity. Regarding precipitation, statistically significant improvements are obtained in equitable threat score, frequency bias and false alarm ratio, particularly in non-summer regimes. Skill enhancements with GDPS-SN are generally found to be more remarkable in winter when large scales apparently play a more critical role in determining weather at the regional level.
In addition to enhancing the skill and predictability of GEM, the proposed hybrid NWP-AI system is shown to generate the full spectra of fine scales that is missing in GC inferences. Furthermore, the new system is capable of producing all the usual forecast variables that operational meteorologists rely on for forecasting weather in general and high-impact events in particular.
@@
Comparisons against near-surface observations have revealed that – unlike GC, which has a limited ability to capture the extremes in wind speed and temperature – weather extremes predicted by GDPS-SN closely follow the operational GDPS-CTL. This implies that the fine scales predicted by GPDS-SN not only possess similar level of energy as GDPS-CTL, they are also comparable in terms of skill.
@@
Other aspects of near-surface predictions from GDPS-SN, however, show relatively modest improvement (e.g., screen-level scores). Lack of vertical resolution in the boundary layer in GC-13 and inconsistencies in surface forcing (such as orography) between ERA5 and GDPS analyses were the reasons why nudging in the boundary layer has not been considered in this study. This largely explains the less-than-remarkable near-surface performance of GDPS-SN. However, work is currently in progress at ECCC to fine-tune GC to emulate the operational GDPS analyses. This fine-tuned version is expected to enhance GC’s skill with the 37-level version, particularly in the stratosphere and the boundary layer. Any such improvement could be easily leveraged by the hybrid system in the near future. Importantly, a fine-tuned GC would allow for nudging in the boundary layer which is expected to result in further enhancement of near-surface skill with GDPS-SN.
@@
Finally, results from the proposed hybrid NWP-AI system imply that, rather than viewing AI and NWP models as two competing paradigms, it is likely more prudent to consider these forecasting methods as complementary. A well-designed fusion of these two approaches can significantly mitigate their individual limitations while allowing for harnessing their respective strengths to provide better meteorological guidance. Therefore, instead of focusing on efforts to replace one with the other, future research should prioritize improving both NWP and AI models.
@@ Acknowledgements.
Figures 20a and 20b illustrate the response of this DCT-based filter for a configuration based on $\lambda_{LS}=1500$ km and $\lambda_{SS}=1000$ km. The figures demonstrate that the DCT-based filter is capable of targeting the desired scales for the individual Yin-Yang sub-domains. The response of the filter over the globe was determined by inspecting the spectral variance ratio of the filtered and unfiltered fields over a global Gaussian grid with respect to a spherical harmonics-based decomposition. The corresponding results are presented in Fig. 20c. The figure reveals that, even though DCT is primarily designed for LAMs where curvature effects of spherical surfaces are neglected, its application on global Yin-Yang grid does not result in any major issue as the filter is found to result in an acceptable response.

### v2

@@ 6.6 Computational cost of spectral nudging
The current implementation of spectral nudging in GEM with the presented optimal configuration leads to an increase of computational cost by approximately 25%. Applying nudging to specific humidity increases the computational cost by another 10% approximately. Due to the proof-of-concept nature of this study, no optimization of the informatics code has been performed. Therefore, the cost could be considerably reduced through future code optimization for the spectral filter and implementation of an asynchronous input server. Further improvements could be attained by considering time-varying nudging increments, where nudging would only be applied at time-steps close to the inference times of GC. These considerations however lie beyond the scope of this study.
@@ 7 Summary and Future Work
The recent emergence of AI models has begun to disrupt the operational weather forecasting paradigm which has long been dominated by the physics-based NWP models. Naturally, serious questions about the future role of AI and NWP models in operational forecasting are arising. For instance, are AI models capable of adding any reliable and meaningful value for weather services, or is it possible to improve guidance from operational NWP models by leveraging the AI models? This study attempts to address these pressing questions by choosing the GEM and GC models to represent the two forecasting approaches.
@@
First, by comparing and contrasting the GEM and GC models, their relative strengths and weaknesses are identified. Skill verification against observations and GDPS analyses reveals that GC suffers from excessive fine-scale smoothing, affecting scales as large as 2500 km beyond day 3 of a 10-day forecast. However, despite smoothing, large-scale features (wavelength$>$2500 km) in GC exhibit better skill compared to GEM, which is evidenced by increased spectral coherence relative to the GDPS analyses and reduced STDE against radiosonde observations.
@@
Inspired by the improved large-scale skill of GC, a hybrid NWP-AI system – namely, GDPS-SN – has been developed to produce global forecasts where GEM-predicted large scales in horizontal wind and temperature are spectrally nudged toward inferences from GC. This hybrid system is capable of generating real-time forecasts with accuracy that surpasses ECCC’s operational GPDS. For example, RMSE of the 500-hPa geopotential height is reduced by 5-10%, with an overall predictability improvement of 12-18 hours, peaking at day 7 of a 10-day forecast. Notable improvements are also obtained in the predicted precipitation and tropical cyclone trajectories. Furthermore, predictions from GDPS-SN are expected to be more physically consistent than GC as, at its core, GDPS-SN employs an NWP model based on the first principles.
@@
In addition to enhancing the skill and predictability of GEM, the proposed hybrid NWP-AI system is shown to generate the full spectra of fine scales that are missing from GC inferences. Comparisons against near-surface observations have revealed that – unlike GC, which has a limited ability to capture the extremes in wind speed and temperature – weather extremes predicted by GDPS-SN closely follow the operational GDPS-CTL.
Operational NWP models are generally required to track hundreds of prognostic variables while generating forecasts with high vertical resolution and temporal frequency. These capabilities are essential for meteorologists to forecast weather patterns, particularly those relevant for high-impact events. The proposed hybrid system maintains these necessary capabilities, albeit with a reasonable increase in computational cost. In this regard, it is important to recognize that similar capability requirements would substantially diminish the computational cost advantage of the AI models. Additionally, training such AI models would be prohibitively expensive.
@@
Work is currently in progress at ECCC to fine-tune GC to emulate the operational GDPS analyses. This fine-tuned version is expected to enhance GC’s skill with the 37-level version, particularly in the stratosphere and the boundary layer. Any such improvement could be easily leveraged by the hybrid system in the near future. Importantly, a fine-tuned 37-level version of GC would allow for nudging in the boundary layer which is expected to result in further enhancement of near-surface skill with GDPS-SN, which with the current optimal configuration, shows less remarkable improvements over GDPS-CTL.
In the future, GDPS-SN could be utilized as the nudging reference for other downstream higher-resolution deterministic GEM-based operational NWP systems. Additionally, spectral nudging could be investigated for global ensemble prediction using a probabilistic AI model, such as GenCast (Price et al., 2024).
@@
Finally, results from this study imply that, rather than viewing AI and NWP models as two competing paradigms, it is likely more prudent to consider these forecasting methods as complementary. A well-designed fusion of these two approaches can significantly mitigate their individual limitations while allowing for harnessing their respective strengths to provide better meteorological guidance. Therefore, instead of focusing on efforts to replace one with the other, future research should prioritize improving both NWP and AI models.
@@ Acknowledgements.
Figures 19a and 19b illustrate the response of this DCT-based filter for the optimal configuration based on $\lambda_{LS}=2750$ km and $\lambda_{SS}=2250$ km. The figures demonstrate that the DCT-based filter is capable of targeting the desired scales for the individual Yin-Yang sub-domains. The response of the filter over the globe was determined by inspecting the spectral variance ratio of the filtered and unfiltered fields over a global Gaussian grid with respect to a spherical harmonics-based decomposition. The corresponding results are presented in Fig. 19c. The figure reveals that, even though DCT is primarily designed for LAMs where curvature effects of spherical surfaces are neglected, its application on global Yin-Yang grid does not result in any major issue as the filter is found to result in an acceptable response.